Country ABC imposes an individual income tax based on the following (progressive) schedule.
5%                                                                                     $0-$50,000

8%                                                                                     $50,001-$200,000

12%                                                                                  $201,000 - ABOVE

Mrs. Jones has $70,000 taxable income. Compute the tax owed on this income.

$19,300

$4,100

$3,500

$5,600
Answer #1

As per the question, for the first $50000, tax at the rate of 5% is to be charged and on the additional $20000 (70000-50000), tax at the rate of 8% is to be charged. Hence, tax owed by him would be (50000*0.05) + (20000*0.08) = (2500+1600) = $4100